OpenStudy (anonymous):

What is the sum of the solutions of 2[x-1]-4=-2?

Directrix (directrix):

Are these brackets for absolute value or the greatest integer function or something else? @kaitlyn9695

OpenStudy (anonymous):

Honestly, I'm not sure. I got this question wrong and I'm trying to figure out how to solve it:/

OpenStudy (anonymous):

This is what it looks like on my paper

OpenStudy (anonymous):

|x-1| means the absolute value of x-1

OpenStudy (anonymous):

Absolute value of 1 = 1 Absolute value of -1 = 1

OpenStudy (anonymous):

Absolute value just means distance away from zero

OpenStudy (anonymous):

so it would be 2 [1]-4=-2

OpenStudy (anonymous):

don't take it into account at first

OpenStudy (anonymous):

solve for x

Directrix (directrix):

>>so it would be 2 [1]-4=-2 No

OpenStudy (shaik0124):

2[x-1]-4=-2? put x=0 2[0-1]-4=2*(-1)-4 =-2-4=-6 since -6 is not equal to -2, 0 is not a solution now try with another number lets take x=2 now 2[2-1]-4=2*1-4 =-2 since -2 is equal to -2 x=2 is a solution similarly consider other values of x which satisfies the equation and add all those which satifies the equation

OpenStudy (shaik0124):

@kaitlyn9695 got it?

OpenStudy (anonymous):

Yes thank you very much! Seeing it broken down was very helpful! This is my last class before I graduate and its a big pain in the butt!

OpenStudy (anonymous):

and the add 2 to each side came from the 4?

OpenStudy (shaik0124):

on the right hand side 2 is negative so -2+2 will be 0

OpenStudy (anonymous):

okay I understand now. When I do future equations like this I just need to plug in the multiple choice option and see if its equal

OpenStudy (shaik0124):

yeah

OpenStudy (shaik0124):

or u can do like this 2[x-1]-4=-2 add 2 on both sides then u will get 2[x-1]-4+2=-2+2 2[x-1]-2=0 again add 2 on both sides then u will get 2[x-1]=2 divide by 2 the equation will be x-1=1 add 1 on both sides x=2

OpenStudy (shaik0124):

Just plugin

OpenStudy (shaik0124):

have u understood the way i solved

OpenStudy (anonymous):

It makes a lot more sense now. Thank you! You're a great teacher

OpenStudy (shaik0124):

u can plugin or directly obtain the value of variable by adding or subtracting (whatever is required

OpenStudy (shaik0124):

U can take my help anytime pleasure to help u

OpenStudy (anonymous):

Thank you, I will let you know if I get stuck again.

OpenStudy (shaik0124):

sure

zepdrix (zepdrix):

\[\large\rm 2|0-1|-4=-2\]Absolute value of -1 is 1,\[\large\rm 2(1)-4=-2\]\[\large\rm -2=-2\] @shaik0124 Is true, so x=0 `is` a solution. They wanted the `sum of the solutions` though, 2 + 0 = 2, so missing that doesn't really affect your answer :)